# Composite Flow Matching for Reinforcement Learning with Shifted-Dynamics Data

### v2

@@ 3.2 Data Collection at High Dynamics Gap Region
As discussed in Section 2.1, using behavior cloning and augmenting the replay buffer with low dynamics gap source data can alleviate performance drop from distribution shift. However, relying solely on such data may limit state-action coverage and hinder policy learning. To address this, we propose a new data collection strategy.
At each training iteration, we construct the replay buffer by selectively incorporating source transitions with small estimated dynamics gap:
=$\mathcal{B}=\left\{(s,a)\in\mathcal{D}_{\text{src}}:\widehat{\Delta}(s,a)\leq\tau\right\}\cup\mathcal{D}_{\text{tar}},$ (3)
@@
where $\beta$ is a hyperparameter that trades off return and exploration of underexplored dynamics.
@@
Compared to behavior cloning on the offline source dataset, training a policy $\hat{\pi}$ by replacing all source samples with a dynamics gap exceeding $\kappa$ (as estimated by the composite flow) with target domain samples can reduce the performance gap to the optimal target policy $\pi^{*}_{\text{tar}}$ by
$\Delta_{\kappa}=\frac{4L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}\left(\Delta_{W_{2}}-\kappa-\sqrt{\text{Var}(s^{\prime}|\nu)_{\text{tar}}+O\left(\frac{\operatorname{Tr}({\Sigma_{s^{\prime}}^{\text{tar}}})}{\sqrt{N_{\text{tar}}}}\right)}\right)$ (5)
Here $L_{r}$ and $L_{P}$ are the Lipschitz constants for the reward and transition function, respectively. $\gamma L_{p}<1$. $\Delta_{W_{2}}\;:=\;\sup_{s,a}W_{2}\!\bigl{(}p_{\text{src}}(\cdot|s,a),\;p_{\text{tar}}(\cdot|s,a)\bigr{)}$ is the largest dynamics gap. $\text{Var}(s^{\prime}|s,a)_{\text{tar}}$ is the conditional variance of $s^{\prime}$ given $s,a$ under the distribution $p_{\text{tar}}$.
From Theorem 3.3, exploration in regions with high dynamics gap can reduce the performance gap relative to the optimal policy. The parameter $\kappa$ serves as a threshold: the bound holds when the policy is trained without source samples whose dynamics gap exceeds $\kappa$. As $\beta$ increases, more samples are collected from high-gap regions, which decreases $\kappa$. This increases the bound $\Delta_{\kappa}$ on performance gap reduction, resulting in a policy with higher expected return.
@@ 3.3 Practical Implementation
Our method can be instantiated using standard actor-critic algorithms with a critic $Q_{\varsigma}(s,a)$ and a policy $\pi_{\varphi}(a|s)$. To incorporate the dynamics gap, we apply rejection sampling to retain a fixed percentage of source transitions with the lowest estimated gap in each iteration. The critic is trained by minimizing
$\mathcal{L}_{Q}=\mathbb{E}_{(s,a,s^{\prime})\sim D_{\text{tar}}}[(Q_{\varsigma}(s,a)-y)^{2}]+\mathbb{E}_{(s,a,s^{\prime})\sim D_{\text{src}}}[\mathbf{1}(\widehat{\Delta}(s,a)\leq\widehat{\Delta}_{\xi\%})(Q_{\varsigma}(s,a)-y)^{2}],$ (6)
where $\widehat{\Delta}_{\xi\%}$ is the $\xi$-quantile of the estimated dynamics gap in the source minibatch. The target value is $y=r+\gamma Q_{\varsigma}(s^{\prime},a^{\prime})+\beta\widehat{\Delta}(s,a),$ with $a^{\prime}\sim\pi_{\varphi}(\cdot|s^{\prime}).$
@@
$\mathcal{L}_{\pi}=\mathbb{E}_{s\sim D_{\rm src}\cup D_{\rm tar},\,a\sim\pi_{\varphi}(\cdot|s)}[Q_{\varsigma}(s,a)]-\omega\,\mathbb{E}_{(s,a)\sim D_{\rm src},\,\tilde{a}\sim\pi_{\varphi}(\cdot|s)}[\|a-\tilde{a}\|^{2}],$ (7)
where $\omega$ is a hyperparameter that balances policy improvement with imitation of source actions. The behavior cloning term encourages sampled actions from the policy to match those in the offline source dataset as suggested in Lemma 2.2.
The pseudocode of CompFlow, instantiated with Soft Actor-Critic (SAC) [18], is presented in Appendix C.
@@ 4 Related Work
Online RL with Offline Dataset. Online RL often requires extensive environment interactions [48, 62], which can be costly or impractical in real-world settings. To improve sample efficiency, Offline-to-Online RL leverages pre-collected offline data to bootstrap online learning [38]. A typical two-phase approach trains an initial policy offline, then fine-tunes it online [38, 28, 39]. However, conservative strategies used to mitigate distributional shift—such as pessimistic value estimation [27]—can produce suboptimal initial policies and hinder exploration [35, 39]. To resolve this tension, recent methods propose ensemble-based pessimism [28], value calibration [39], optimistic action selection [28], and policy expansion [64]. Others directly incorporate offline data into the replay buffer of off-policy algorithms [3], improving stability via ensemble distillation and layer normalization. However, these works assume identical dynamics between the source and target environments. In contrast, our work explicitly addresses the dynamics shift between the offline and online environments.
@@
Domain Adaptation for RL. Our setting aligns with domain adaptation, where the source and target domains share observation and action spaces but differ in transition dynamics. Prior work has addressed such discrepancies via system identification [8, 10, 5], domain randomization [50, 55, 36], imitation learning [23, 19], and meta-RL [37, 45], often assuming shared environment distributions [52] or requiring expert demonstrations. Recent methods emphasize dynamics-aware policy learning. One direction is reward modification, which penalizes source transitions unlikely under target dynamics by adjusting the reward function [12, 30] or down-weighting value estimates in high-gap regions [41]. Another is data filtering, which retains only source transitions with low estimated dynamics gap [59], using metrics such as transition probability ratios [12], mutual information [58], value inconsistency [59], or representation KL divergence [32]. However, these metrics can be ill-defined when dynamics have disjoint support, and value-based methods suffer from instability due to bootstrapping bias.
@@
Our work differs from these approaches by leveraging the theoretical connection between optimal transport and flow matching, which enables principled estimation of the Wasserstein distance between source and target transition as the dynamics gap. A closely related work by [34] also employs optimal transport to quantify the dynamics gap. However, their setting assumes both the source and target domains are offline, and their estimation is based on the concatenated tuple $(s,a,s^{\prime})$ observed in the offline dataset, rather than computing a distance between the conditional transition distributions $p_{\rm src}(s^{\prime}|s,a)$ and $p_{\rm tar}(s^{\prime}|s,a)$. Consequently, their metric does not provide an accurate measure of the transition dynamics gap and cannot be directly used to compute exploration bonuses, which require dynamics gap estimation conditioned on a given state-action pair.
@@
RL with Diffusion and Flow Models. Diffusion models [20, 51] and flow matching [29, 11] have emerged as powerful generative tools capable of modeling complex, high-dimensional distributions. Their application in RL is consequently expanding. Researchers have employed these generative models for various tasks, including planning and trajectory synthesis [21], representing expressive multimodal policies [7, 46], providing behavior regularization [6], or augmenting training datasets with synthesized experiences. While these works often focus on policy learning or modeling dynamics within a single environment, our approach targets the transfer learning setting. Specifically, we address scenarios characterized by a dynamics gap between source and target domains. We utilize Flow Matching, leveraging its connection to optimal transport, to estimate this gap.
@@ 5 Experiments
In this section, we evaluate our approach across a variety of environments featuring different types of dynamics shifts. We also perform ablation and hyperparameter studies to gain deeper insights into the design and behavior of CompFlow.
5.1 Experimental Setup
@@
Tasks and datasets. We evaluate our algorithm under two types of dynamics shifts—morphology and kinematic—across three OpenAI Gym locomotion tasks: HalfCheetah, Hopper, and Walker2d [4]. Each experiment involves an offline source domain and an online target domain with modified transition dynamics following [33]. Morphology shifts change body part sizes, while kinematic shifts constrain joint angles. For each task, we use three D4RL source datasets—medium, medium-replay, and medium-expert—capturing varying data quality [16]. Additional results and environment details are in Appendix I and J.
@@
Baselines. We compare CompFlow against the following baselines: BC-SAC extends SAC by incorporating both source and target data, with a behavior cloning (BC) term for the offline source. H2O [41] penalizes Q-values for state-action pairs with large dynamics gaps. BC-VGDF [59] selects source transitions with value targets consistent with the target domain and adds a BC term. BC-PAR [32] applies a reward penalty based on representation mismatch between source and target transitions, also including a BC term. Implementation details are provided in Appendix J.
5.2 Main Results
Dataset Task Name BC-SAC H2O BC-VGDF BC-PAR Ours
MR HalfCheetah (Morphology) $2495\pm 43$ $1430\pm 408$ $2765\pm 124$ $1790\pm 91$ $3119\pm 107$
MR HalfCheetah (Kinematic) $4868\pm 186$ $4257\pm 609$ $4392\pm 403$ $4179\pm 441$ $5189\pm 262$
MR Hopper (Morphology) $346\pm 4$ $361\pm 18$ $348\pm 21$ $354\pm 25$ $355\pm 6$
MR Hopper (Kinematic) $1024\pm 0$ $1025\pm 0$ $1024\pm 0$ $1024\pm 1$ $1024\pm 1$
MR Walker2D (Morphology) $598\pm 475$ $1014\pm 193$ $672\pm 576$ $458\pm 151$ $1094\pm 791$
M HalfCheetah (Morphology) $1522\pm 72$ $1720\pm 273$ $1829\pm 345$ $1427\pm 196$ $2282\pm 287$
M HalfCheetah (Kinematic) $5451\pm 195$ $5019\pm 773$ $4972\pm 381$ $5243\pm 120$ $5593\pm 44$
M Hopper (Morphology) $436\pm 45$ $410\pm 8$ $406\pm 52$ $418\pm 13$ $604\pm 173$
M Hopper (Kinematic) $1022\pm 1$ $970\pm 98$ $934\pm 43$ $1020\pm 3$ $1023\pm 2$
M Walker2D (Morphology) $457\pm 317$ $577\pm 201$ $584\pm 219$ $431\pm 177$ $886\pm 372$
ME HalfCheetah (Morphology) $1195\pm 241$ $1147\pm 169$ $1072\pm 102$ $1207\pm 53$ $1485\pm 67$
ME HalfCheetah (Kinematic) $4211\pm 262$ $5143\pm 330$ $4603\pm 498$ $4399\pm 164$ $5750\pm 84$
ME Hopper (Morphology) $349\pm 47$ $444\pm 15$ $357\pm 63$ $407\pm 28$ $462\pm 89$
ME Hopper (Kinematic) $1024\pm 1$ $1031\pm 3$ $1022\pm 3$ $1027\pm 8$ $1022\pm 2$
ME Walker2D (Morphology) $429\pm 117$ $1103\pm 444$ $502\pm 301$ $380\pm 231$ $648\pm 180$
Average Return $1695$ $1710$ $1699$ $1584$ $2036$

### v3

@@ 3.2 Data Collection at High Dynamics Gap Region
As discussed in Section 2.1, using behavior cloning and augmenting the replay buffer with low dynamics gap offline data can alleviate performance drop from distribution shift. However, relying solely on such data may limit state-action coverage and hinder policy learning. To address this, we propose a new data collection strategy.
At each training iteration, we construct the replay buffer by selectively incorporating offline transitions with small estimated dynamics gap:
$\mathcal{B}=\left\{(s,a)\in\mathcal{D}_{\text{off}}:\widehat{\Delta}(s,a)\leq\tau\right\}\cup\mathcal{D}_{\text{on}},$ (4)
@@
where $\beta$ is a hyperparameter that trades off return and exploration of underexplored dynamics. To ensure sufficient coverage during training, we can further incorporate stochasticity by adding small perturbations to the selected actions [fujimoto2018addressing] or following a stochastic policy [haarnoja2018soft] consistent with widely used deep RL frameworks.
@@
Compared to behavior cloning policy $\pi_{\rm bc}$ on the offline dataset, training a policy $\hat{\pi}$ by replacing all offline samples with a dynamics gap exceeding $\kappa$ (as estimated by the composite flow) with online environment samples can reduce the performance gap to the optimal online policy $\pi^{*}_{\text{on}}$ with high probability by
$\frac{2L_{r}(1+\gamma)}{(1-\gamma)(1-\gamma L_{p})}\bigl(\Delta_{W_{2}}-\kappa-\sqrt{\left(C_{0}+C_{1}\,W_{2}(\hat{p}_{\rm off},p_{\rm on})\right)\Gamma_{N_{\rm on},\delta}}\bigr).$ (6)
Here $L_{r}$ and $L_{P}$ are the Lipschitz constants for the reward and transition function, respectively. $\gamma L_{p}<1$. $\Delta_{W_{2}}\;:=\;\sup_{s,a}W_{2}\!\bigl(p_{\rm off}(\cdot|s,a),\;p_{\rm on}(\cdot|s,a)\bigr)$ is the largest dynamics gap. $C_{0}$ and $C_{1}$ are two constants. $\Gamma_{N_{\rm on},\delta}:=\mathfrak{R}_{N_{\rm on}}(\mathcal{H})\;+\;\sqrt{\tfrac{\log(1/\delta)}{N_{\rm on}}}$, where $\mathcal{H}$ is the same as in Theorem 3.1 and $N_{\mathrm{on}}$ is the number of samples used to train the online flow.
From Theorem 3.4, exploration in regions with high dynamics gap can reduce the performance gap relative to the optimal policy. The parameter $\kappa$ serves as a threshold: the bound holds when the policy is trained without offline samples whose dynamics gap exceeds $\kappa$. As $\beta$ increases, more samples are collected from high-gap regions, which decreases $\kappa$. This increases performance gap reduction, resulting in a policy with higher expected return.
@@ 3.3 Practical Implementation
Our method can be instantiated using standard actor-critic algorithms with a critic $Q_{\varsigma}(s,a)$ and a policy $\pi_{\varphi}(a|s)$. To incorporate the dynamics gap, we apply rejection sampling to retain a fixed percentage of offline transitions with the lowest estimated gap in each iteration. The critic is trained by minimizing
$\mathcal{L}_{Q}=\mathbb{E}_{(s,a,s^{\prime})\sim D_{\text{on}}}[(Q_{\varsigma}(s,a)-y)^{2}]+\mathbb{E}_{(s,a,s^{\prime})\sim D_{\text{off}}}[\mathbf{1}(\widehat{\Delta}(s,a)\leq\widehat{\Delta}_{\xi\%})(Q_{\varsigma}(s,a)-y)^{2}],$ (7)
where $\widehat{\Delta}_{\xi\%}$ is the $\xi$-quantile of the estimated dynamics gap in the offline minibatch. The target value is $y=r+\gamma Q_{\varsigma}(s^{\prime},a^{\prime})+\beta\widehat{\Delta}(s,a),$ with $a^{\prime}\sim\pi_{\varphi}(\cdot|s^{\prime}).$
@@
$\mathcal{L}_{\pi}=\mathbb{E}_{s\sim D_{\rm off}\cup\mathcal{D}_{\rm on},\,a\sim\pi_{\varphi}(\cdot|s)}[Q_{\varsigma}(s,a)]-\omega\,\mathbb{E}_{(s,a)\sim\mathcal{D}_{\rm off},\,\tilde{a}\sim\pi_{\varphi}(\cdot|s)}[\|a-\tilde{a}\|^{2}],$ (8)
where $\omega$ is a hyperparameter that balances policy improvement with imitation of source actions. The behavior cloning term encourages sampled actions from the policy to match those in the offline dataset as suggested in Lemma 2.2.
The pseudocode of CompFlow, instantiated with Soft Actor-Critic (SAC) [haarnoja2018soft], is presented in Appendix D.
@@ 4 Related Work
Online RL with offline dataset. Online RL often requires extensive environment interactions [silver2016mastering, ye2020towards], which can be costly or impractical in real-world settings. To improve sample efficiency, Offline-to-Online RL leverages pre-collected offline data to bootstrap online learning [nair2020awac]. A typical two-phase approach trains an initial policy offline, then fine-tunes it online [nair2020awac, lee2022offline, nakamoto2023cal]. However, conservative strategies used to mitigate distributional shift, such as pessimistic value estimation [kumar2020conservative], can result in suboptimal initial policies and limit effective exploration [mark2022fine, nakamoto2023cal]. To resolve this tension, recent methods propose ensemble-based pessimism [lee2022offline], value calibration [nakamoto2023cal], optimistic action selection [lee2022offline], and policy expansion [zhang2023policy]. Others directly incorporate offline data into the replay buffer of off-policy algorithms [ball2023efficient], improving stability via ensemble distillation and layer normalization. However, these works assume identical dynamics between the source and target environments. In contrast, our work explicitly addresses the dynamics shift between the offline and online environments.
@@
RL with dynamics shift. Our work is related to cross-domain RL, where the source and target domains share the same observation and action spaces but differ in transition dynamics. Prior work has addressed such discrepancies via system identification [clavera2018learning, du2021auto, chebotar2019closing], domain randomization [slaoui2019robust, tobin2017domain, mehta2020active], imitation learning [kim2020domain, hejna2020hierarchically], and meta-RL [nagabandi2018learning, raileanu2020fast], often assuming shared environment distributions [song2020provably] or requiring expert demonstrations. More recent work has relaxed these assumptions. One line of research focuses on reward modification, which adjusts the reward function to penalize source transitions that are unlikely under the target dynamics [eysenbach2020off, liu2022dara], or down-weights value estimates in regions with high dynamics gap [niu2022trust]. Another line of work explores data filtering, which selects only source transitions with low estimated dynamics gap [xu2023cross], using metrics such as transition probability ratios [eysenbach2020off], mutual information [wen2024contrastive], value inconsistency [xu2023cross], or representation-based KL divergence [lyu2024cross]. However, these metrics can become unstable or ill-defined when the transition dynamics have different supports, and value-based methods are prone to instability caused by bootstrapping bias.
@@
In contrast, our approach leverages the theoretical connection between optimal transport and flow matching to estimate the dynamics gap in a principled manner. A closely related work by [lyu2025cross] also applies optimal transport to quantify the dynamics gap. However, their setting assumes both the source and target domains are offline, and their method estimates the gap based on the concatenated tuple $(s,a,s^{\prime})$ observed in offline data, rather than comparing conditional transition distributions. As a result, their metric does not accurately capture the gap in transition dynamics and cannot be used to compute exploration bonuses, which require gap estimation conditioned on a given state-action pair.
@@
RL with diffusion and flow models. Diffusion models [ho2020denoising, song2020score, kong2024diffusion] and flow matching [lipman2023flow, esser2024scaling] have emerged as powerful generative tools capable of modeling complex, high-dimensional distributions. Their application in RL is consequently expanding. Researchers have employed these generative models for various tasks, including planning and trajectory synthesis [janner2022planning], representing expressive multimodal policies [chi2023diffusionpolicy, ren2025diffusion], providing behavior regularization [chen2024score], or augmenting training datasets with synthesized experiences. While these works often focus on policy learning or modeling dynamics within a single environment, our approach targets the transfer learning setting. Specifically, we address scenarios characterized by a dynamics gap between the offline data and the online environment. We utilize Flow Matching, leveraging its connection to optimal transport, to estimate this gap.
@@ 5 Experiments
In this section [^1] Our code is available at https://github.com/Haichuan23/CompositeFlow, we first evaluate our approach across a range of environments in Gym-MuJoCo that exhibit different types of dynamics shifts. We then conduct ablation and hyperparameter studies to better understand the design choices and behavior of CompFlow. Finally, we assess the effectiveness of our method in a real-world inspired wildlife conservation task.
5.1 Gym-MuJoCo
5.1.1 Experimental Setup
@@
Tasks and datasets. We evaluate our algorithm under three types of dynamics shifts, namely morphology, kinematic and friction changes, across three OpenAI Gym locomotion tasks: HalfCheetah, Hopper, and Walker2d [brockman2016openai]. Each experiment involves an offline environment and an online environment with modified transition dynamics following [lyu2024odrl]. Morphology shifts alter the sizes of body parts, kinematic shifts impose constraints on joint angles, and friction shifts modify the static, dynamic, and rolling friction coefficients. For each task, we use three D4RL source datasets: medium, medium replay, and medium expert, which capture varying levels of data quality [fu2020d4rl]. Additional environment details are in Appendix K.
@@
Baselines. We compare CompFlow against the following baselines: BC-SAC extends SAC by incorporating both offline and online data, with a behavior cloning (BC) term for the offline data. H2O [niu2022trust] penalizes Q-values for state-action pairs with large dynamics gaps. BC-VGDF [xu2023cross] selects offline transitions with value targets consistent with the online environment and adds a BC term. BC-PAR [lyu2024cross] applies a reward penalty based on representation mismatch between offline and online transitions, also including a BC term. Implementation details are provided in Appendix K.
5.1.2 Main Results
Dataset Task Name SAC BC-SAC H2O BC-VGDF BC-PAR Ours
MR HalfCheetah (Morphology) $1457\pm 89$ $2495\pm 43$ $1430\pm 408$ $2765\pm 124$ $1790\pm 91$ $3119\pm 107$
MR HalfCheetah (Kinematic) $2255\pm 197$ $4868\pm 186$ $4257\pm 609$ $4392\pm 403$ $4179\pm 441$ $5189\pm 262$
MR HalfCheetah (Friction) $2069\pm 184$ $7799\pm 157$ $6397\pm 673$ $7829\pm 821$ $8056\pm 512$ $8241\pm 180$
MR Hopper (Morphology) $364\pm 82$ $346\pm 4$ $361\pm 18$ $348\pm 21$ $354\pm 25$ $355\pm 6$
MR Hopper (Kinematic) $737\pm 547$ $1024\pm 0$ $1025\pm 0$ $1024\pm 0$ $1024\pm 1$ $1024\pm 1$
MR Hopper (Friction) $34\pm 4 22$$8\pm 2 22$$29\pm 1$ $230\pm 3$ $232\pm 5$ $280\pm 27$
MR Walker2D (Morphology) $253\pm 60$ $598\pm 475$ $1014\pm 193$ $672\pm 576$ $458\pm 151$ $1094\pm 791$
MR Walker2D (Kinematic) $152\pm 22$ $2973\pm 185$ $1967\pm 851$ $1586\pm 923$ $948\pm 131$ $1568\pm 1315$
MR Walker2D (Friction) $301\pm 9$ $311\pm 5$ $296\pm 14$ $302\pm 12$ $321\pm 26$ $344\pm 20$
M HalfCheetah (Morphology) $1467\pm 89$ $1522\pm 72$ $1720\pm 273$ $1829\pm 345$ $1427\pm 196$ $2282\pm 287$
M HalfCheetah (Kinematic) $2316\pm 92$ $5451\pm 195$ $5019\pm 773$ $4972\pm 381$ $243\pm 120 55$$593\pm 44$
M HalfCheetah (Friction) $2028\pm 238$ $7108\pm 1001$ $6968\pm 846$ $6802\pm 956$ $7800\pm 525$ $7871\pm 238$
M Hopper (Morphology) $396\pm 60$ $436\pm 45$ $410\pm 8$ $406\pm 52$ $418\pm 13$ $604\pm 173$
M Hopper (Kinematic) $724\pm 535$ $1022\pm 1$ $970\pm 98$ $934\pm 43$ $1020\pm 3$ $1023\pm 2$
M Hopper (Friction) $229\pm 5$ $32\pm 1 22$$8\pm 4 22$$29\pm 4$ $233\pm 5$ $300\pm 66$
M Walker2D (Morphology) $301\pm 177$ $457\pm 317$ $577\pm 201$ $584\pm 219$ $431\pm 177$ $886\pm 372$
M Walker2D (Kinematic) $258\pm 174$ $1966\pm 1155$ $1965\pm 568$ $1921\pm 928$ $806\pm 278$ $2039\pm 936$
M Walker2D (Friction) $301\pm 9$ $286\pm 54$ $298\pm 45$ $289\pm 47$ $308\pm 24$ $320\pm 31$
ME HalfCheetah (Morphology) $392\pm 238 11$$95\pm 241 11$$147\pm 169$ $1072\pm 102$ $1207\pm 53$ $1485\pm 67$
ME HalfCheetah (Kinematic) $2323\pm 97$ $4211\pm 262$ $5143\pm 330$ $4603\pm 498$ $4399\pm 164$ $5750\pm 84$
ME HalfCheetah (Friction) $1950\pm 312$ $4185\pm 732$ $2140\pm 733$ $4078\pm 1032$ $4989\pm 500$ $5596\pm 1557$
ME Hopper (Morphology) $359\pm 75$ $349\pm 47$ $444\pm 15$ $357\pm 63$ $407\pm 28$ $462\pm 89$
ME Hopper (Kinematic) $724\pm 535$ $1024\pm 1$ $1031\pm 3$ $1022\pm 3$ $1027\pm 8$ $1022\pm 2$
ME Hopper (Friction) $229\pm 4$ $230\pm 3$ $232\pm 5$ $230\pm 6$ $232\pm 8$ $266\pm 70$
ME Walker2D (Morphology) $228\pm 51$ $429\pm 117$ $1103\pm 444$ $502\pm 301$ $380\pm 231$ $648\pm 180$
ME Walker2D (Kinematic) $386\pm 184$ $850\pm 953$ $1514\pm 782$ $1204\pm 734$ $755\pm 268$ $1511\pm 1206$
ME Walker2D (Friction) $266\pm 66$ $240\pm 114$ $258\pm 8$ $245\pm 51$ $242\pm 24$ $326\pm 26$
Average Return $\mathbf{878}$ $\mathbf{1920}$ $\mathbf{1783}$ $\mathbf{1868}$ $\mathbf{1803}$ $\mathbf{2193}$